The p.d.f of a continuous random variable is give


P(x) ={ kx (1-x) e*, 0 ≤ x ≤ 1

{ 0, otherwise

Find k and hence find mean and standard deviation



1
Expert's answer
2021-06-27T17:26:05-0400
"1=\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}kx(1-x)dx"

"=k\big[\dfrac{x^2}{2}-\dfrac{x^3}{3}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{k}{6}"

"k=6"

"f(x)=\begin{cases}
 6x(1-x), & 0\leq x\leq 1 \\
 0, &\text{otherwise} 
\end{cases}"


"\mu=E[X]=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{1}6xx(1-x)dx"

"=6\big[\dfrac{x^3}{3}-\dfrac{x^4}{4}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=2-\dfrac{3}{2}=\dfrac{1}{2}=0.5"



"E[X^2]=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{1}6x^2x(1-x)dx"

"=6\big[\dfrac{x^4}{4}-\dfrac{x^5}{5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{3}{2}-\dfrac{6}{5}=\dfrac{3}{10}"




"V(X)=\sigma^2=E[X^2]-(E[X])^2"

"=\dfrac{3}{10}-(\dfrac{1}{2})^2=\dfrac{1}{20}=0.05"

"\sigma=\sqrt{\sigma^2}=\sqrt{0.05}=0.1\sqrt{5}\approx0.2236"